OpenStudy (anonymous):

Solve this equation plz: 7^2+x=78

OpenStudy (anonymous):

7*7 = 49 78 - 49 = ? X = the answer to 78 -49

OpenStudy (mwhit2101):

7^2 is 49 so 49+x=78 Subtract 49 from both sides X equals 29

OpenStudy (anonymous):

THank you guys very much

OpenStudy (anonymous):

Your welcome @MackenzieGirl